OpenStudy (anonymous):

Short = 5 , hypotenuse = 10 Long =

OpenStudy (anonymous):

\[(short)^2+(long)^2=(hypotenuse)^2\]

OpenStudy (anonymous):

Have you read about right triangles?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):



OpenStudy (anonymous):

yes been there but havent seen the problem lay out before

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

\[short=a\] \[long=b\] \[hyp=c\]\[a^2+b^2=c^2\]\[b^2=c^2-a^2\] can you find the b(long) now?